Fermi National Laboratory


Accelerator Update for March 16, 2005 - March 18, 2005

Wednesday March 16
The day shift began with Operations monitoring store 4044 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE, NuMI, and Meson MCenter experiments taking beam.

A Main Injector (MI) expert began a slipstacking study at 9:52 AM.

A NuMI expert began a quadrupole study at 9:56 AM.

Booster experts began tuning their beam at 3 PM.

Pbar experts completed a reverse proton tune-up and then began a study period at 3:18 PM. Stacking stopped for the duration of the study.

NuMI studies ended at 4:25 PM and they resumed normal running.

Pbar studies ended at 8:45 PM and Pbar resumed stacking.

Operations terminated store 4044 at 9:04 PM.

Operators prepared the Collider Detector at Fermilab (CDF) and D-Zero (D0) for access at 9:25 PM. Both experiments were out of access by 11:05 PM.

At 11:31 PM, Operations reported that the CDF turn on was being held up due to a quench protection monitor (QPM) problem. An expert said a corrupt file was the cause.

Thursday March 17
The midnight shift began with Operations monitoring the stack and the stash, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations put the Tevatron (TeV) through a dry squeeze at 12:08 AM.

Operators turned off NuMI and MCenter in preparation for shot setup at 12:41 AM.

Operations established store 4046 at 4:28 AM with an initial luminosity of 98.18E30. Pbar resumed stacking. The eighth transfer of antiprotons was lost due to a Pbar sequencer problem.

NuMI and MCenter resumed taking beam at 5:24 AM.

The Neutron Therapy Facility (NTF) requested beam at 10:21 AM to tune their beam.

At 10:33 AM, Operations turned off MI to allow an access into the MI-31 stub.

Operation made the Linac ready for access 10:37 AM.

Preaccelerator experts removed the I- Source at 10:47 AM. They had it replaced within an hour.

Around 11 AM, a TeV expert began a helix study.

Operations established beam in the Linac at 11:54 AM.

Operations established beam to Pbar, NuMI, and MCenter at 11:57 AM.

Operations put the Recycler in to shot setup to stash antiprotons at 2:06 PM.

Recycler shots ended and Pbar resumed stacking at 3:26 PM.

The TeV helix study ended around 3:45 PM.

D0 experimenters reported at 5:32 PM that they were testing their luminosity counters.

Operations established beam to the Meson MTest experiment at 5:38 PM.

At 5:39 PM, Operations reported that a MI RF station (MIRF18) tripped off for the first of four times in a little less than two hours. It reset okay each time.

At 5:57 PM, Operations reported that MIRF9 tripped off for the first of three times in about an hour. It reset okay each time.

MTest experimenters reported at 8:05 PM that they were finished with the beam for the night.

Friday March 18
The midnight shift began with Operations monitoring store 4046 and a small stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 4:54 AM, a TeV RF station (TRF6) tripped off. The trip didn't seem to affect CDF, but D0 took an antiproton and luminosity hit.

At 5:54 AM, a D0 cold compressor came into alarm. Later, D0 called and said their wet engine had tripped off.

The Plan for Friday and the Weekend
The plan for today is to terminate the store around 9 AM. The Run Coordinator will allow CDF and D0 to make an hour access after the store is terminated. Operations will put the TeV into shot setup around 11 AM. The rest of the day and the weekend will consist of stacking, stashing, stores, and increasing intensity for NuMI.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/18/2005   email Fermilab